# Gated Tree Cross-attention for Checkpoint-Compatible Syntax Injection in Decoder-Only LLMs

### v1

@@ Abstract
Decoder-only large language models achieve strong broad performance but are brittle to minor grammatical perturbations, undermining reliability for downstream reasoning. However, directly injecting explicit syntactic structure into an existing checkpoint can interfere with its pretrained competence. We introduce a checkpoint-compatible gated tree cross-attention (GTCA) branch that reads precomputed constituency chunk memory while leaving backbone architecture unchanged. Our design uses a token update mask and staged training to control the scope and timing of structural updates. Across benchmarks and Transformer backbones, GTCA strengthens syntactic robustness beyond continued-training baselines without compromising Multiple-Choice QA performance or commonsense reasoning, providing a practical checkpoint-compatible route to more syntax-robust decoder-only LLMs.
@@
Xinyu Gao1  Shaonan Wang2  Nai Ding1∗ 1Zhejiang University 2The Hong Kong Polytechnic University
@@ 1 Introduction
Evaluation covers multiple-choice QA (MCQA) benchmarks (CLOTH (Xie et al., 2018), MMLU (Hendrycks et al., 2021)), syntactic benchmarks (BLiMP (Warstadt et al., 2020), CoLA Warstadt et al. (2019) from GLUE (Wang et al., 2018)) and common sense reasoning tasks (HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021)). GTCA consistently improves syntax-focused results while maintaining MCQA accuracy and overall capability across different backbones. Layer-wise probe of Unlabeled Undirected Attachment Score (UUAS) further connects these gains to a more syntax-consistent internal structure (Hewitt and Manning, 2019; Müller-Eberstein et al., 2022; Limisiewicz and Mareček, 2021). Specifically, GTCA increases the accuracy of BLiMP from 78.58 to 83.12 in Qwen-2.5-7B and from 79.95 to 84.61 in Llama-3-8B, despite leaving the backbone architecture unchanged.
@@ 2.1 Injecting explicit syntactic structure into pre-trained Transformers
Another direction injected syntax through training objectives rather than architectural rewrites, e.g., syntax-guided contrastive learning (Zhang et al., 2022), or continuing training pipelines that explicitly considered optimization stability and catastrophic forgetting when incorporating syntactic knowledge (Iwamoto et al., 2023). In parallel, modeling native language structure on a scale aims to induce or enforce hierarchical structure during pretraining (Hu et al., 2024; Zhao et al., 2024), and graph-based formulations offer a broader perspective of structured modeling (Plenz and Frank, 2024). At the same time, cautionary evidence suggests that tree bias alone is not sufficient. Its effectiveness depends on how the structure is represented and integrated (Ginn, 2024). These findings motivate checkpoint-compatible, minimally invasive structure injection for decoder-only LLMs, where hierarchical signals are exposed through an auxiliary pathway and integrated in a way that the pre-trained backbone can reliably exploit.
@@ 2.2 Probing syntactic representations and training-efficient adaptation
Probing has been central to characterizing syntactic information in neural language models and to clarifying the mechanisms-outcomes tension. Structural probing showed that dependency-like geometry could be recovered from contextual embeddings (Hewitt and Manning, 2019), while attention and layer-wise analyzes have revealed systematic associations between model components and linguistic relations or processing stages (Jawahar et al., 2019; Tenney et al., 2019; Clark et al., 2019). Subsequent work improved probe formulations and constraints (Limisiewicz and Mareček, 2021; Müller-Eberstein et al., 2022), extended probing to incremental parse states in autoregressive settings (Eisape et al., 2022), and proposed richer representational codes such as polar probes (Diego-Simón et al., 2024).
However, recent evidence has emphasized that representational recoverability does not guaranty behavioral usage. Redundancy and probe artifacts can produce misleading conclusions, motivating intervention-style and dropout-based analyzes (Tucker et al., 2022; Agarwal et al., 2025; Diego-Simón et al., 2025), while other work argued for a richer generative-style structure than previous probes captured (Kennedy, 2025). In parallel, practical checkpoint adaptation has been shaped by parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2021) and QLoRA (Dettmers et al., 2023), prompt-based alternatives including prefix and prompt tuning (Li and Liang, 2021; Lester et al., 2021; Liu et al., 2022), and stability-oriented components such as gated attention variants that improve optimization dynamics (Qiu et al., 2025). Together, this literature motivates approaches that control interference during continued training and that evaluate syntax gains through both targeted behavioral tests and layer-wise structural probes. These principles are reflected in our staged training pipeline and our joint evaluation.
Figure 1: Overview of GTCA (Gated Tree Cross-Attention) structural injection.Given a input (question, options, and answer field), we run an offline parser with span alignment and encode the resulting hierarchy with a parse-tree encoder to build per-layer chunk memory $C^{\ell}$. The chunk memory is injected into a frozen backbone layer via gated cross-attention, and the resulting structural update is used to update token states $H^{\ell}$.
@@ 3.1 Precomputed parse tree structure aligned to token spans
Concretely, we define a binary token update mask $mask_{\text{token}}\in\{0,1\}^{n}$ that enables structural updates only for question tokens and the answer field, while forcing $mask_{\text{token}}=0$ for option tokens. This prevents option preference drift during continued training and confines structural intervention to the parts of the input where syntactic reasoning is intended to operate.
@@ 3.2 Hierarchical parse tree encoder for chunk memory construction
$\tilde{\ell}=\min(\ell,D),\quad C_{\ell}=\left[c_{u}\right]_{u\in U,\,h(u)=\tilde{\ell}}$ (3)
@@
This layer-aligned design, where the chunk $C^{\ell}$ is consumed by the Transformer layer $\ell$, is intended to encourage an intuitive hierarchy of structural utilization, such that the higher layers draw on the higher-level parse chunks while the lower layers primarily rely on more local structure. We empirically probe this effect via UUAS (Section 4.4).
@@ 3.3 Gated parse tree cross-attention injected via wrappers
We inject parse tree structure through an auxiliary cross-attention branch implemented as a forward wrapper, using head-wise gated attention (Qiu et al., 2025). The headwise gate is adopted because previous work showed that it introduces far fewer parameters than the elementwise gate while achieving comparable perplexity and downstream accuracy, thus reducing memory overhead and stabilizing training. At Transformer layer $\ell$, let $C^{\ell}$ denote the chunk memory at height $\ell$ (defined in Section 3.2), and $\overline{H^{\ell}}$ denote the pre-normalization token hidden states. We use projection matrices $W_{Q}^{\ell},W_{K}^{\ell},W_{V}^{\ell}$ and $W_{G}^{\ell}$, then compute queries, keys, values, gate score, and attention output with head dimension $d_{h}$:
$\begin{split}Q^{\ell}&=\overline{H^{\ell}}W_{Q}^{\ell},K^{\ell}=C^{\ell}W_{K}^{\ell},\\ V^{\ell}&=C^{\ell}W_{V}^{\ell},G^{\ell}=\overline{H^{\ell}}W_{G}^{\ell}\end{split}$ (4)
@@
$H^{\ell}\leftarrow H^{\ell}+\alpha_{\text{struct}}\bigl(mask_{\text{token}}[:,\text{None}]\odot\Delta H^{\ell}\bigr)$ (8)
where the token update mask $mask_{\text{token}}\in\{0,1\}^{n}$ is defined in Section 3.1 and the structural coefficient $\alpha_{\text{struct}}$ follows the schedule in Section 3.4. Here the token update mask $mask_{\text{token}}[:,\text{None}]$ indicates that it is broadcast along the hidden dimension so that each row of $H^{\ell}$ is multiplied by the corresponding mask entry.
@@ 4.1 Experimental Setup
Unless otherwise stated, all experiments in the main manuscript are reported in both Qwen-2.5-7B (Qwen et al., 2025) and Llama-3-8B (Grattafiori et al., 2024), using the same maximum sequence length and prompts, computing budgets, and the number of training steps, while keeping the default tokenizer of each backbone. Additionally, Appendix A provides parameter counts for the 8 variants (4 variants × 2 backbones), along with dataset details, the overall compute budget and computing infrastructure.
@@
For structure aware variants, we precomputed constituency parse trees offline, cached the resulting parse tree memory. To ensure that all examples could be successfully parsed, we applied a minimal text normalization to the raw data by collapsing consecutive whitespace into a single space, and modeled each sentence independently in all the experiments. For MCQA prompts, we parsed each field independently (question, each option, and the answer prefix), rather than parsing the full concatenated prompt. We then concatenated the resulting span-aligned chunks in the original prompt order. In addition, Appendix B provides the prompts used for each benchmark along with the corresponding dataset, checkpoint licensing and terms of use.
@@
Table 3: Total scores on CLOTH, BLiMP, MMLU, COLA, HellaSwag and Winogrande datasets for the Backbone, LoRA-only, Direct-Joint, and GTCA.
Training details.
We followed the three stages training scheme in Section 3.4. The key hyperparameters are reported in Table 2. We selected the structural coefficient $\alpha_{\text{struct}}^{*}=0.15$ through grid search (Appendix C).
Evaluation details.
$\mathrm{score}\!\left(option^{j}\mid prompt\right)=$ (9)
where $j\in\{1,\ldots,r\}$ indexes candidate options and $i$ indexes tokens within $option^{j}$. And our token update mask could keep the option log-likelihoods in Equation LABEL:eqq comparable across candidates, thereby cleanly reflecting the influence of syntactic structure in the tree.
The prediction was the option with the highest score:
@@ 4.3 Retention, robustness, and syntactic generalization
In the general capability benchmarks HellaSwag and Winogrande, GTCA delivers strong general capacity retention across both backbones. In both Qwen-2.5-7B and Llama-3-8B, GTCA achieves the best HellaSwag performance among the continued-training baselines. For Winogrande, GTCA matches LoRA-only in Qwen-2.5-7B and is strongest in Llama-3-8B. Although GTCA is not uniformly best on every general benchmark and backbone, it achieves top performance in most settings and remains competitive with the strongest baselines elsewhere. Together, these results indicate that staged training and gated integration produce measurable syntactic gains while maintaining broad general capability largely stable.
@@ 4.4 Layer-wise syntactic probing
Table 4: Ablation results on key model components. No Mask sets $mask_{\text{token}}=1$ for all tokens.
@@
Table 5: Performance dynamics across specialization stages.
5 Analysis
We analyzed which components drove the BLiMP and CoLA gains of syntax injection, and how they interacted with retention on general benchmarks. We first ablated gating, token update masking, and staged training, then tested whether improvements depended on coherent parse structure rather than generic auxiliary capacity. To isolate component effects without introducing confounding differences from backbone and tokenizer-specific differences, we conducted these follow-up analyzes on Qwen-2.5-7B, while the main results already demonstrate consistent trends across both backbones.
5.1 Ablations and stage-wise effects
Table 4 presents controlled ablations of the entire model. Overall, GTCA consistently delivers the best trade-off between syntactic generalization and retention, suggesting that each design choice contributes meaningfully rather than acting as redundant extra capacity. In particular, removing the gate degrades both the targeted syntax in BLiMP and the broad competence in MMLU, highlighting the role of the gate in controlling interference from the structural pathway. In contrast, disabling the token update mask leaves syntactic performance essentially unchanged but slightly weakens MCQA-format stability and general retention, indicating that token update masking mainly stabilized continued training under the MCQA interface rather than directly driving syntactic gains.
Table 5 makes the stage-wise dynamics explicit. Stage 2 delivers the greatest improvement in BLiMP (about +2.6 points), showing that freezing the backbone and specializing structural modules is where syntactic generalization is most strongly acquired. This gain comes with a temporary drop in MMLU, but Stage 3 largely restores and even improves broad competence while preserving most of the syntactic improvement. This pattern suggests that Stage 3 mitigates the retention drop introduced in Stage 2 while preserving most syntactic gains.
@@ 5.1 Ablations and stage-wise effects
Table 6: Tree quality and structure-faithfulness controls. Strong parser* corresponds to GTCA (full parser).

### v2

@@ Abstract
Decoder-only large language models achieve strong broad performance but are brittle to minor grammatical perturbations, undermining reliability for downstream reasoning. However, directly injecting explicit syntactic structure into an existing checkpoint can interfere with its pretrained competence. We introduce a checkpoint-compatible gated tree cross-attention (GTCA) branch that reads precomputed constituency chunk memory while leaving backbone architecture unchanged. Our design uses a token update mask and staged training to control the scope and timing of structural updates. Across benchmarks and Transformer backbones, GTCA strengthens syntactic robustness beyond continued training baselines without compromising Multiple-Choice QA performance or commonsense reasoning, providing a practical checkpoint-compatible route to more syntax-robust decoder-only LLMs [^1] https://github.com/Pineandgrass/GatedTreeCrossAttention..
@@
Xinyu Gao1, Shaonan Wang2, Nai Ding1thanks: Corresponding author: Nai Ding 1College of Biomedical Engineering & Instrument Science, Zhejiang University 2Department of Language Science and Technology, The Hong Kong Polytechnic University {xinyu_gao, ding_nai}@zju.edu.cn shaonan.wang@polyu.edu.hk
@@ 1 Introduction
Evaluation covers multiple-choice QA (MCQA) benchmarks, including CLOTH (Xie et al., 2018) and MMLU (Hendrycks et al., 2021), syntactic benchmarks, including BLiMP (Warstadt et al., 2020) and CoLA Warstadt et al. (2019) from GLUE (Wang et al., 2018), and commonsense reasoning tasks, including HellaSwag (Zellers et al., 2019) and WinoGrande (Sakaguchi et al., 2021). GTCA consistently improves syntax-focused results while maintaining MCQA accuracy and overall capability across different backbones. Layer-wise probe of Unlabeled Undirected Attachment Score (UUAS) further connects these gains to a more syntax-consistent internal structure (Hewitt and Manning, 2019; Müller-Eberstein et al., 2022; Limisiewicz and Mareček, 2021). Specifically, GTCA increases the accuracy of BLiMP from 78.58 to 83.12 in Qwen-2.5-7B and from 79.95 to 84.61 in Llama-3-8B, despite leaving the backbone architecture unchanged.
@@ 2.1 Injecting Explicit Syntactic Structure into Pre-Trained Transformers
A related line of work injected syntax through training objectives rather than architectural rewrites, e.g., syntax-guided contrastive learning (Zhang et al., 2022), or continuing training pipelines that explicitly considered optimization stability and catastrophic forgetting when incorporating syntactic knowledge (Iwamoto et al., 2023). In parallel, modeling native language structure on a scale aims to induce or enforce hierarchical structure during pretraining (Hu et al., 2024; Zhao et al., 2024), and graph-based formulations offer a broader perspective of structured modeling (Plenz and Frank, 2024). At the same time, cautionary evidence suggests that tree bias alone is not sufficient. Its effectiveness depends on how the structure is represented and integrated (Ginn, 2024). These findings motivate checkpoint-compatible, minimally invasive structure injection for decoder-only LLMs, where hierarchical signals are exposed through an auxiliary pathway and integrated in a way that the pre-trained backbone can reliably exploit.
@@ 2.2 Probing Syntactic Representations and Training-Efficient Adaptation
Probing has been central to characterizing syntactic information in neural language models and to clarifying the mechanisms-outcomes tension. Structural probing showed that dependency-like geometry could be recovered from contextual embeddings (Hewitt and Manning, 2019), while attention and layer-wise analyses have revealed systematic associations between model components and linguistic relations or processing stages (Jawahar et al., 2019; Tenney et al., 2019; Clark et al., 2019). Subsequent work improved probe formulations and constraints (Limisiewicz and Mareček, 2021; Müller-Eberstein et al., 2022), extended probing to incremental parse states in autoregressive settings (Eisape et al., 2022), and proposed richer representational codes such as polar probes (Diego-Simón et al., 2024).
However, recent evidence has emphasized that representational recoverability does not guarantee behavioral usage. Redundancy and probe artifacts can produce misleading conclusions, motivating intervention-style and dropout-based analyses (Tucker et al., 2022; Agarwal et al., 2025; Diego-Simón et al., 2025), while other work argued for a richer generative-style structure than previous probes captured (Kennedy, 2025). In parallel, practical checkpoint adaptation has been shaped by parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2021) and QLoRA (Dettmers et al., 2023), prompt-based alternatives including prefix and prompt tuning (Li and Liang, 2021; Lester et al., 2021; Liu et al., 2022), and stability-oriented components such as gated attention variants that improve optimization dynamics (Qiu et al., 2025). Together, this literature motivates approaches that control interference during continued training and that evaluate syntax gains through both targeted behavioral tests and layer-wise structural probes. These principles are reflected in our staged training pipeline and our joint evaluation.
Figure 1: Overview of GTCA (Gated Tree Cross-Attention) structural injection.Given an input (question, options, and answer field), we run an offline parser with span alignment and encode the resulting hierarchy with a parse-tree encoder to build per-layer chunk memory $C^{\ell}$. At Transformer layer $\ell$, the pre-update token states $H_{\mathrm{pre}}^{\ell}$ query $C^{\ell}$ through gated tree cross-attention to produce a structural residual update $\Delta H^{\ell}$. This update is then added to obtain the post-update token states $H_{\mathrm{post}}^{\ell}$, which are passed to the next layer.
@@ 3.1 Precomputed Parse Tree Structure Aligned to Token Spans
Concretely, we define a binary token update mask $m_{\mathrm{tok}}\in\{0,1\}^{n}$ that enables structural updates only for question tokens and the answer field, while forcing $m_{\mathrm{tok}}=0$ for option tokens. This prevents option preference drift during continued training and confines structural intervention to the parts of the input where syntactic reasoning is intended to operate.
@@ 3.2 Hierarchical Parse Tree Encoder for Chunk Memory Construction
$\tilde{\ell}=\min(\ell,D),\quad C^{\ell}=\left[c_{u}\right]_{u\in U,\,h(u)=\tilde{\ell}}$ (3)
@@
This layer-aligned design, where the chunk $C^{\ell}$ is consumed by the Transformer layer $\ell$, is intended to encourage an intuitive hierarchy of structural utilization, such that the higher layers draw on the higher-level parse chunks while the lower layers primarily rely on more local structure. We empirically probe this effect via UUAS (Section 4.4). Controlled ablations on Qwen-2.5-7B for the projection scheme, span pooling, and memory size are reported in Appendix A.
@@ 3.3 Gated Parse Tree Cross-Attention Injected via Wrappers
We inject parse tree structure through an auxiliary cross-attention branch implemented as a forward wrapper, using head-wise gated attention (Qiu et al., 2025). The headwise gate is adopted because previous work showed that it introduces far fewer parameters than the elementwise gate while achieving comparable perplexity and downstream accuracy, thus reducing memory overhead and stabilizing training.
At Transformer layer $\ell$, let $C^{\ell}$ denote the chunk memory at height $\ell$ (defined in Section 3.2), and $\overline{H^{\ell}_{\mathrm{pre}}}$ denote the pre-normalization token hidden states before the GTCA update. We use projection matrices $W_{Q}^{\ell},W_{K}^{\ell},W_{V}^{\ell}$ and $W_{G}^{\ell}$, then compute queries, keys, values, gate score, and attention output with head dimension $d_{h}$:
$\begin{split}Q^{\ell}&=\overline{H^{\ell}_{\mathrm{pre}}}W_{Q}^{\ell},\quad K^{\ell}=C^{\ell}W_{K}^{\ell},\\ V^{\ell}&=C^{\ell}W_{V}^{\ell},\quad G^{\ell}=\overline{H^{\ell}_{\mathrm{pre}}}W_{G}^{\ell},\end{split}$ (4)
@@
$H^{\ell}_{\mathrm{post}}\leftarrow H^{\ell}_{\mathrm{pre}}+\alpha_{\text{struct}}\bigl(m_{\mathrm{tok}}[:,\text{None}]\odot\Delta H^{\ell}\bigr)$ (8)
where the token update mask $m_{\mathrm{tok}}\in\{0,1\}^{n}$ is defined in Section 3.1 and the structural coefficient $\alpha_{\text{struct}}$ follows the schedule in Section 3.4. Here $m_{\mathrm{tok}}[:,\mathrm{None}]$ denotes broadcasting along the hidden dimension, so that each row of $\Delta H^{\ell}$ is multiplied by the corresponding mask entry before the masked update is added to $H_{\mathrm{pre}}^{\ell}$ to obtain $H_{\mathrm{post}}^{\ell}$.
@@ 4.1 Experimental Setup
Unless otherwise stated, all experiments in the main manuscript are reported in both Qwen-2.5-7B (Qwen et al., 2025) and Llama-3-8B (Grattafiori et al., 2024), using the same maximum sequence length and prompts, computing budgets, and the number of training steps, while keeping the default tokenizer of each backbone. Additionally, Appendix B provides parameter counts for the 8 variants (4 variants × 2 backbones), along with dataset details, the overall compute budget and computing infrastructure.
@@
For structure-aware variants, we precomputed constituency parse trees offline, cached the resulting parse tree memory. To ensure that all examples could be successfully parsed, we applied a minimal text normalization to the raw data by collapsing consecutive whitespace into a single space, and modeled each sentence independently in all the experiments. For MCQA prompts, we parsed each field independently (question, each option, and the answer prefix), rather than parsing the full concatenated prompt. We then concatenated the resulting span-aligned chunks in the original prompt order. In addition, Appendix C provides the prompts used for each benchmark along with the corresponding dataset, checkpoint licensing and terms of use.
@@
Training details.
We followed the three-stage training scheme in Section 3.4. The key hyperparameters are reported in Table 2. We selected the structural coefficient $\alpha_{\text{struct}}^{*}=0.15$ through grid search (Appendix D).
Due to the checkpoint-compatible design, the absolute differences between the backbone parameters before and after training are 0.0 for both backbones, indicating that the original backbone checkpoint parameters remain unchanged.
Evaluation details.
$\mathrm{score}\!\left(option^{j}\mid prompt\right)=$ (9)
where $j\in\{1,\ldots,r\}$ indexes candidate options and $i$ indexes tokens within $option^{j}$. And our token update mask can keep the option log-likelihoods in Equation LABEL:eqq comparable across candidates, thereby cleanly reflecting the influence of syntactic structure in the tree. The prediction was the option with the highest score:
When option lengths vary strongly, length normalized option scoring is preferable, but in our setting it leaves the method ranking unchanged.
@@
Table 3: Total scores on CLOTH, BLiMP, MMLU, CoLA, HellaSwag, and WinoGrande for Backbone, LoRA-only, Direct-Joint, and GTCA. Boldface with $\dagger$ indicates the best result when the improvement is significant under paired tests with Holm–Bonferroni correction ($p<0.05$ after correction).
@@ 4.3 Retention, Robustness, and Syntactic Generalization
In the general capability benchmarks HellaSwag and WinoGrande, GTCA delivers strong general capacity retention across both backbones. In both Qwen-2.5-7B and Llama-3-8B, GTCA achieves the best HellaSwag performance among the continued training baselines. For WinoGrande, GTCA matches LoRA-only in Qwen-2.5-7B and is strongest in Llama-3-8B. Although GTCA is not uniformly best on every general benchmark and backbone, it achieves top performance in most settings and remains competitive with the strongest baselines elsewhere. Together, these results indicate that staged training and gated integration produce measurable syntactic gains while maintaining broad general capability largely stable.
@@ 4.4 Layer-Wise Syntactic Probing
5 Analysis
We analyzed which components drove the BLiMP and CoLA gains of syntax injection, and how they interacted with retention on general benchmarks. We first ablated gating, token update masking, and staged training, then tested whether improvements depended on coherent parse structure rather than generic auxiliary capacity. To isolate component effects without introducing confounding differences from backbone and tokenizer-specific differences, we conducted these follow-up analyses on Qwen-2.5-7B, while the main results already demonstrate consistent trends across both backbones.
5.1 Ablations and Stage-Wise Effects
Table 4 presents controlled ablations of the entire model. Overall, GTCA consistently delivers the best trade-off between syntactic generalization and retention, suggesting that each design choice contributes meaningfully rather than acting as redundant extra capacity. In particular, removing the gate degrades both the targeted syntax in BLiMP and the broad competence in MMLU, highlighting the role of the gate in controlling interference from the structural pathway. In contrast, disabling the token update mask leaves syntactic performance essentially unchanged but slightly weakens MCQA format stability and general retention. On a held-out set in Qwen-2.5-7B, the mean absolute change in pairwise option score gaps is 0.013 per token with the mask and 0.021 without it, consistent with the No Mask ablation in Table 4. This suggests that token update masking mainly stabilizes continued training under the MCQA interface rather than directly driving syntactic gains.
Table 5 makes the stage-wise dynamics explicit and also includes a two stage pipeline that skips Stage 2. Stage 2 delivers the greatest improvement in BLiMP (about +2.6 points), showing that freezing the backbone and specializing structural modules is where syntactic generalization is most strongly acquired. This gain comes with a temporary drop in MMLU, but Stage 3 largely restores and even improves broad competence while preserving most of the syntactic improvement. This pattern suggests that Stage 2 and Stage 3 play complementary roles and that skipping either one yields a weaker tradeoff.
We further conducted a qualitative error analysis on BLiMP in Appendix E, where the remaining errors are concentrated in filler gap dependencies and island effects, suggesting that the model can still over-rely on local surface compatibility and underweight global constraints that require tracking long-distance dependencies or respecting extraction islands.
@@ 5.1 Ablations and Stage-Wise Effects
Table 4: Ablation results on key model components. No Mask sets $m_{\mathrm{tok}}=1$ for all tokens. Boldface with $\dagger$ indicates the best result when the improvement is significant under paired tests with Holm–Bonferroni correction ($p<0.05$ after correction).
@@
Table 5: Performance dynamics across specialization stages. Boldface with $\dagger$ indicates the best result when the improvement is significant under paired tests with Holm–Bonferroni correction ($p<0.05$ after correction).
@@
Table 6: Tree quality and structure-faithfulness controls. Strong parser* corresponds to GTCA (full parser). Boldface with $\dagger$ indicates the best result when the improvement is significant under paired tests with Holm–Bonferroni correction ($p<0.05$ after correction).